OpenStudy (anonymous):

Create a quadratic equation with these two solutions. 8, 0

OpenStudy (anonymous):

solution is always \[(x-r_1)(x-r_2)...(x-r_n)=0\] where the \[r_i\] are the solutions. here you have \[r_1=0,r_2=8\] \[(x-0)(x-8)=0\] \[x(x-8)=0\] \[x^2-8x=0\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome